OpenStudy (anonymous):

Simplify the sum. 7/(a+8) + 7/(a^2 - 64)

OpenStudy (anonymous):

Okay so notice that the bottom of the second fraction is infact the "difference of two squares" i.e. (a^2-64)=(x-8)(x+8), then multiply the first fraction by x-8 so that you have equal denominators then you may add them together simply and simplify if nessecary

OpenStudy (ajprincess):

What is the lcm of a+8 and a^2-64?

OpenStudy (anonymous):

64?

OpenStudy (lacypennelll):

You play AnimalJam?

OpenStudy (ajprincess):

a+8=1*(a+8) a^2-64=(a+8)(a-8)

OpenStudy (lacypennelll):

.-. Thats wha aj stands for

OpenStudy (anonymous):

aj??

OpenStudy (lacypennelll):

^ nvm

OpenStudy (anonymous):

aj u there?

OpenStudy (ajprincess):

yup I am

OpenStudy (anonymous):

oh ok so that answer i will is C or D

OpenStudy (anonymous):

??

OpenStudy (lacypennelll):

:| Silence

OpenStudy (ajprincess):

Lcm of a+8 and a^2-64 is a^2-64 a^2-64=(a-8)(a+8) So the common denominator is a^2-64 With vat expression do u need to multiply (a+8) to make it a^2-64?

OpenStudy (ajprincess):